OpenStudy (anonymous):

Please help me :'( Write x^2 - 6x + 8= 0 in vertex form. Write your answer in the form a(x-h)^2+k=0

OpenStudy (triciaal):

general form ax^2 + bx + c a = 1 ;b = -6 ; c = 8 vertex form a(x-h)^2 + k a = a ; h = -b/2a; k = c + h^2

OpenStudy (triciaal):

correction k = c - h^2

OpenStudy (arabpride):

Please close this post if you're done. Thanks! :D